Reports 1-1 of 1 Clear search Modify search
Optical characterization (Optical characterization)
magazzu, nardecchia, gouaty - 19:28 Tuesday 07 May 2024 (64206) Print this report
Scan of the OMC in CARM_NULL_1F

Figure 1 shows the scan of the OMC performed from 16h08 utc to 16h30 utc, after waiting for 1 hour in CARM_NULL_1F.

Figure 2 shows a zoom of this scan around a Free Spectral Range.

Assuming a calibration factor of about 4800 for B1_PD3, we have: ~75 mW on the first order mode, ~29 mW on the second order mode, ~200 mW on each of the third and fourth order modes, ~70 mW on each of the 5th and 6th order modes, ...


Images attached to this report
Comments to this report:
mwas - 12:31 Monday 13 May 2024 (64238) Print this report

Figure 1 tries to fit the scan calibrated into MHz, the scan is not well linear so it doesn't work perfectly well. The order 2 mode is crossed at least 4 times during the scan, these are the 4 peaks highlighted, two during the time the temperature is increasing and two during the time the temperature is decreasing. The order 2 mode varies between 19mW and 41mW using the same calibration factor for B1 PD3 of 4800.


Images attached to this comment
Search Help